After seeing an alert at 11:20, network administrators discovered a failed management module in the 2nd floor switch. The module was swapped and the configuration reloaded. The switch was returned to service at 11:53am. Instructional support staff had to reboot several servers which had become stuck due to disrupted NFS traffic. All 2nd floor Instructional Labs were operational as of 12:30.
Resolved as of 2009-09-23 11:53:00